# GOMP: Grasped Object Manifold Projection for Multimodal Imitation Learning of Manipulation

### v1

@@ Abstract
Imitation Learning (IL) holds great potential for learning repetitive manipulation tasks, such as those in industrial assembly. However, its effectiveness is often limited by insufficient trajectory precision due to compounding errors. In this paper, we introduce Grasped Object Manifold Projection (GOMP), an interactive method that mitigates these errors by constraining a non-rigidly grasped object to a lower-dimensional manifold. GOMP assumes a precise task in which a manipulator holds an object that may shift within the grasp in an observable manner and must be mated with a grounded part. Crucially, all GOMP enhancements are learned from the same expert dataset used to train the base IL policy, and are adjusted with an $n$-arm bandit-based interactive component. We propose a theoretical basis for GOMP’s improvement upon the well-known compounding error bound in IL literature. We demonstrate the framework on four precise assembly tasks using tactile feedback, and note that the approach remains modality-agnostic. Data and videos are available at williamvdb.github.io/GOMPsite.
@@ I Introduction
To address this challenge, we introduce Grasped Object Manifold Projection (GOMP), an interactive framework which operates on top of an IL policy to constrain a grasped object to a lower dimensional task space, as outlined in Fig 1. Our method learns a task-space manifold from expert demonstrations and projects IL trajectories to this manifold, removing compounding errors orthogonal to its tangent space. We also provide an interactive reinforcement learning method for selecting the manifold based on observed successes. Finally, we demonstrate GOMP against vanilla IL—implemented as a Diffusion Policy [chi2023]—on real robot experiments.
@@ II Related Works
GOMP does not rely on dataset aggregation as DAgger does. Rather, it assumes an existing, constant, task-specific manipulation dataset, and applies extra understanding of a task manifold to the behavior-cloned policy. This is inspired by several Bayesian and reward-based methods. DropoutDAgger [ken2017] ensures the predicted action is sufficiently close to the expert action and invokes the expert if not. SQIL [red2019] provides simple rewards for novice behavior that is closer to the expert. CCIL [ke2024_ccil] also maintains closeness to the expert by generating corrective labels for training. These methods serve a similar purpose to the projection onto the task manifold learned from expert demonstration as in GOMP. GOMP bases its adjustment of this manifold on classical RL, which is used in conjunction with IL in other methods. Inverse Reinforcement Learning (IRL) is a classic method which learns a reward function for RL from demonstration, while Generative Adversarial Imitation Learning (GAIL) [ho2016] extracts a policy directly from demonstrations as if IRL was used.
The state of the art for “vanilla” behavior cloning (imitation learning via learning the state-action map) is Diffusion Policy (DP) [chi2023], where the state-action map is a generative diffusion model. This method has been shown to require >>100 demonstrations for manipulation tasks that require medium precision [chi2023]. For GOMP, we show an improvement upon DP for high precision assembly tasks. In this paper, we demonstrate this with tactile feedback. While DP was initially only tested using visual feedback, it lends itself well to multimodal feedback and has since been tested with tactile feedback, including vision-based tactile feedback. While [geo2024] found DP with tactile feedback to produce zero successful USB insertions, we find some success, with greater success when GOMP is used.
@@ III-B Background and Motivation
This formulation implies that compounding errors accumulate across all states [ross2010]. However, many tasks, particularly those encountered in manufacturing evolve over lower-dimensional manifolds in the state space. These manifolds are induced by the mechanical constraints of a task (e.g., screw motion for nut and bolt assembly or slide-to-insert for peg-insertion). Inspired by this insight, we propose Grasped Object Manifold Projection (GOMP), which constrains the object to a lower-dimensional task space manifold $\mathcal{T}$ during IL rollout. Our method results in a combined policy $\hat{\pi}_{\mathbf{s}_{o}}$ that provides corrections to the base IL policy via projections onto this manifold. Given that the manifold $\mathcal{T}$ is derived from observations of the task-specific expert policy $\pi^{*}$, we assume $\mathcal{T}$ is representative of the task completed by $\pi^{*}$. Thus, we expect the compounding error represented by Eq. 2 are mitigated along directions orthogonal to $\mathcal{T}$.
@@ IV Method
In this section, we describe our instantiation of GOMP, first describing the projection which is applied to the IL trajectories, then outlining the selection process of the projection dimensionality from the expert demonstrations, and the $n$-arm bandit method for adjusting the projection selection from the resulting success rates.
@@ IV-A Task Manifold Projection
Our goal is to derive a behavior cloned policy using the formulation from Ross et al. (Eq. 1). GOMP supplements this problem with the mapping $\mathbf{F}_{\mathbf{s}_{o}}:\mathrm{SE}(3)\rightarrow\mathrm{SE}(3)$ which is computed from a lower-dimensional task space manifold $\mathcal{T}\subseteq\mathrm{SE}(3)$ that is representative of the task. This policy seeks the robot pose $\mathbf{T}_{st}$ such that the object pose is projected to the task space $\mathcal{T}$ which is learned from the demonstration dataset using the method described in Section IV-B. We formalize this computation as:
@@ V Experimental Setup
In this section, we describe the way in which we tested GOMP using a tactile and proprioceptive representation of $\mathbf{s}_{r}$, and a tactile-derived in-hand pose estimator such that $\mathbf{T}_{to}$ is measurable in accordance with our assumptions stated in Section IV-A.
@@ V-B2 Tactile In-Hand Pose Estimation
To recover $\mathbf{T}_{to}$ for GOMP, we use an in-hand pose estimation module $\mathbf{IHP}$, consisting of a convolutional variational auto-encoder [vae] and MLP, with layers as listed in Table I. These are trained simultaneously on reconstruction of the raw RGB GelSlim image (downsampled 16x, 6-channel from two fingers), minimization of KL-divergence of the latent space, and regression to ground truth object poses. This is encompassed in the loss $\mathcal{L}_{\text{ihp}}$:
@@ V-C Observation Encoding
Encoder
$\mathbf{IHP}_{\text{Enc}}$
Conv2D In: $6$, Out: $12$, $3\times 3$, MaxPool($2$)
@@
Predictor
$\mathbf{IHP}_{\text{MLP}}$
Linear In: $256$, Out: $84$
@@
Decoder
$\mathbf{IHP}_{\text{Dec}}$
Linear In: $256$, Out: $256$
@@ V-D Task Demonstrations
We collect a tactile and proprioceptive dataset for each task the robot is to perform; both are important to our implementation of both GOMP and IL.
@@ V-D2 Demonstration Processing
All demonstration episodes for behavior cloning a single task are sampled such that each episode contributes $T_{E}=64$ action/observation pairs, where $\mathbf{A}^{0}_{t}$ consists of $T_{a}=8$ cartesian robot poses $\mathbf{T}_{st}$ (converted to 9D pose) within the $\mathrm{SE}(3)$ robot workspace. Using position (as opposed to velocity) as an action space in this way is consistent with [chi2023] and lends itself well to GOMP. Actions are normalized to (1,1)(-1,1) just as observations. We augment the dataset 8 times by adding noise $\sim N(0,0.1)$ to the normalized tactile shear-field. We use a train-validation split ratio of 80:20. For every 10 episodes, we then have 4096 action-observation pairs to train, and 1024 to validate. The weights of $\epsilon_{\theta}$ that produce the lowest $\mathcal{L}_{D}$ during training are selected for testing.
@@ VI Experiments and Results
We tested GOMP as described in Section IV with DP as described in Section V, against pure vanilla DP as a baseline, with four precise assembly tasks: nut threading, peg insertion, USB insertion, and battery cover placement. We treated this as an interactive imitation learning trial where demonstration episodes are added every 10 trials to train new policies. For the DP baseline, we just tested the effect of adding episodes, beginning with 10, then 20, 40, 60, 80, and finally 100. For GOMP we tested the same, with the effect of the projection to $\mathcal{T}$ (Eq. 4). The initial value $Q_{0}(i)$ was determined by the first 10 episodes only. We set the value update in Eq. 8 to occur every $K=1$ trial, with $\gamma=0.025$. The $\epsilon$-greedy method for selecting the projection dimensionality $i_{k}^{*}$ is performed with $\epsilon=0.1$. Each experiment—a full set of 6 different amounts of training episodes, tested 10 times each—is replicated 4 times. For each of these 4 replications, episodes are introduced in a new random order. As a result, each task is tested with 240 runs for GOMP and 240 runs for the DP baseline. Each run is given a maximum policy horizon: $T_{E}=64$ action prediction steps to complete the task. Snapshots of the four tasks we test are found in Figs. 58.
Figure 5: Nut threading results. Top: Handoff, initialization, policy, and success snapshots. Left: GOMP vs DP performance results as demonstrations are added to training. Right: Change of highest values in $Q(i)$ from Section IV-C over the 60-trial horizon, averaged over the 4 runs of GOMP. Filled area surrounding curves represents $\sigma$ (the initial projection loss in Eq. 7 sometimes yielded $i_{k}^{*}=1$ but $i_{k}^{*}=2$ was the eventual result in all 4 runs of USB insertion).
@@ VI-A Nut Threading
Figure 6: Peg insertion results. Top: Handoff, initialization, policy, and success snapshots. Left: GOMP vs DP performance results as demonstrations are added to training. Right: Change of highest values in $Q(i)$ from Section IV-C over the 60-trial horizon, averaged over the 4 runs of GOMP. Filled area surrounding curves represents $\sigma/8$ (high variance in the result of Eq. 9 occurred between the 4 runs of peg insertion).
@@ VI-B Peg Insertion
Figure 7: USB insertion results. Top: Handoff, initialization, policy, and success snapshots. Left: GOMP vs DP performance results as demonstrations are added to training. Right: Change of highest values in $Q(i)$ from Section IV-C over the 60-trial horizon, averaged over the 4 runs of GOMP. Filled area surrounding curves represents $\sigma$ (Eq. 9 always yielded $i_{k}^{*}=2$ in all 4 runs of USB insertion).
@@ VI-C USB Insertion
Figure 8: Battery cover placement results. Top: Handoff, initialization, policy, and success snapshots. Left: GOMP vs DP performance results as demonstrations are added to training. Right: Change of highest values in $Q(i)$ from Section IV-C over the 60-trial horizon, averaged over the 4 runs of GOMP. Filled area surrounding curves represents $\sigma$ (Eq. 9 yielded $i_{k}^{*}=2$ most commonly, with $i_{k}^{*}=3$ being selected near the end of 1 of 4 runs of battery cover placement).
@@ VII Discussion
Our results suggest that GOMP improves task performance over vanilla Diffusion Policy (DP). One key effect of GOMP is the tendency to avoid unrecoverable states in contrast to vanilla IL, especially when the object shifts in the grasp. For example, nut threading with DP alone may exhibit behavior where the nut contacts the bolt such that it tilts heavily within the grasp. GOMP allows the robot to position itself such that the nut remains vertical, even though this robot configuration does not occur in DP’s expert dataset.
GOMP’s behavior is beneficially reactive to disturbances. This same kind of tilting can happen in peg insertion, USB insertion, and battery cover placement as well. A secondary recovery strategy requires obstacle avoidance. For instance, our peg insertion task is tested with a hole that does not have an excess of surface surrounding its mouth, instead there is empty space. If the peg is erroneously plunged into this empty space rather than the hole, a re-lift of the peg is required to recover, but this behavior does not appear in the expert dataset, nor is it likely to occur with GOMP’s projection. GOMP is more likely to prevent this kind of erroneous behavior to begin with, leading to our more successful peg insertion results seen in Fig. 6.
@@ VII-A Anomalous Baseline Performance
The reason for this paradoxical outcome is debatable, though we hypothesize some form of overfitting is occurring, possibly due to suboptimal, repetitive, or noisy demonstrations. There are also several hyperparameter and architecture choices when it comes to Diffusion Policy. We do note that testing over this set is prohibitively expensive given the need for real-world rollouts. Our results show that GOMP successfully improves upon Diffusion Policy for these tasks, independent of the possible existence of suboptimal demonstrations or design choices.
@@ VII-B Limitations
Future improvements may be inspired by GOMP’s limitations. GOMP currently does not learn a temporal manifold, meaning projecting to $\mathcal{T}$ will not necessarily progress the task temporally. DP retains full responsibility for task completion, while GOMP attempts to prevent accumulating errors. GOMP also requires a reliable in-hand pose estimator. Any improvements in the pose estimation method will serve to improve GOMP. For instance, we demonstrated a predictable $\mathrm{SE}(2)$ estimator given our assumptions (Sec. V-B). GOMP can be improved if $\mathbf{IHP}$ were expanded to $\mathrm{SE}(3)$ as some works have attempted  [t3]. We note that tactile may not be available for all systems. A secondary solution would be to use vision techniques to estimate poses; however, these techniques are known to struggle with occlusion.
Our hope is that the inspiration behind GOMP leads to the consideration of more geometry-based methods on top of the benefits that imitation learning already provides. Along this research path, the data efficiency and precision required by manipulation in fixtureless industrial assembly can be fully realized.

### v2

# GrOMP: Grasped Object Manifold Projection for Multimodal Imitation Learning of Manipulation

@@ Abstract
Imitation Learning (IL) holds great potential for learning repetitive manipulation tasks, such as those in industrial assembly. However, its effectiveness is often limited by insufficient trajectory precision due to compounding errors. In this paper, we introduce Grasped Object Manifold Projection (GrOMP), an interactive method that mitigates these errors by constraining a non-rigidly grasped object to a lower-dimensional manifold. GrOMP assumes a precise task in which a manipulator holds an object that may shift within the grasp in an observable manner and must be mated with a grounded part. Crucially, all GrOMP enhancements are learned from the same expert dataset used to train the base IL policy, and are adjusted with an $n$-arm bandit-based interactive component. We propose a theoretical basis for GrOMP’s improvement upon the well-known compounding error bound in IL literature. We demonstrate the framework on four precise assembly tasks using tactile feedback, and note that the approach remains modality-agnostic. Data and videos are available at williamvdb.github.io/GrOMPsite.
@@ I Introduction
To address this challenge, we introduce Grasped Object Manifold Projection (GrOMP), an interactive framework which operates on top of an IL policy to constrain a grasped object to a lower dimensional task space, as outlined in Fig 1. Our method learns a task-space manifold from expert demonstrations and projects IL trajectories to this manifold, removing compounding errors orthogonal to its tangent space. We also provide an interactive reinforcement learning method for selecting the manifold based on observed successes. Finally, we demonstrate GrOMP against vanilla IL—implemented as a Diffusion Policy [chi2023]—on real robot experiments.
@@ II Related Works
GrOMP does not rely on dataset aggregation as DAgger does. Rather, it assumes an existing, constant, task-specific manipulation dataset, and applies extra understanding of a task manifold to the behavior-cloned policy. This is inspired by several Bayesian and reward-based methods. DropoutDAgger [ken2017] ensures the predicted action is sufficiently close to the expert action and invokes the expert if not. SQIL [red2019] provides simple rewards for novice behavior that is closer to the expert. CCIL [ke2024_ccil] also maintains closeness to the expert by generating corrective labels for training. These methods serve a similar purpose to the projection onto the task manifold learned from expert demonstration as in GrOMP. GrOMP bases its adjustment of this manifold on classical RL, which is used in conjunction with IL in other methods. Inverse Reinforcement Learning (IRL) is a classic method which learns a reward function for RL from demonstration, while Generative Adversarial Imitation Learning (GAIL) [ho2016] extracts a policy directly from demonstrations as if IRL was used.
The state of the art for “vanilla” behavior cloning (imitation learning via learning the state-action map) is Diffusion Policy (DP) [chi2023], where the state-action map is a generative diffusion model. This method has been shown to require >>100 demonstrations for manipulation tasks that require medium precision [chi2023]. For GrOMP, we show an improvement upon DP for high precision assembly tasks. In this paper, we demonstrate this with tactile feedback. While DP was initially only tested using visual feedback, it lends itself well to multimodal feedback and has since been tested with tactile feedback, including vision-based tactile feedback. While [geo2024] found DP with tactile feedback to produce zero successful USB insertions, we find some success, with greater success when GrOMP is used.
@@ III-B Background and Motivation
This formulation implies that compounding errors accumulate across all states [ross2010]. However, many tasks, particularly those encountered in manufacturing evolve over lower-dimensional manifolds in the state space. These manifolds are induced by the mechanical constraints of a task (e.g., screw motion for nut and bolt assembly or slide-to-insert for peg-insertion). Inspired by this insight, we propose Grasped Object Manifold Projection (GrOMP), which constrains the object to a lower-dimensional task space manifold $\mathcal{T}$ during IL rollout. Our method results in a combined policy $\hat{\pi}_{\mathbf{s}_{o}}$ that provides corrections to the base IL policy via projections onto this manifold. Given that the manifold $\mathcal{T}$ is derived from observations of the task-specific expert policy $\pi^{*}$, we assume $\mathcal{T}$ is representative of the task completed by $\pi^{*}$. Thus, we expect the compounding error represented by Eq. 2 are mitigated along directions orthogonal to $\mathcal{T}$.
@@ IV Method
In this section, we describe our instantiation of GrOMP, first describing the projection which is applied to the IL trajectories, then outlining the selection process of the projection dimensionality from the expert demonstrations, and the $n$-arm bandit method for adjusting the projection selection from the resulting success rates.
@@ IV-A Task Manifold Projection
Our goal is to derive a behavior cloned policy using the formulation from Ross et al. (Eq. 1). GrOMP supplements this problem with the mapping $\mathbf{F}_{\mathbf{s}_{o}}:\mathrm{SE}(3)\rightarrow\mathrm{SE}(3)$ which is computed from a lower-dimensional task space manifold $\mathcal{T}\subseteq\mathrm{SE}(3)$ that is representative of the task. This policy seeks the robot pose $\mathbf{T}_{st}$ such that the object pose is projected to the task space $\mathcal{T}$ which is learned from the demonstration dataset using the method described in Section IV-B. We formalize this computation as:
@@ V Experimental Setup
In this section, we describe the way in which we tested GrOMP using a tactile and proprioceptive representation of $\mathbf{s}_{r}$, and a tactile-derived in-hand pose estimator such that $\mathbf{T}_{to}$ is measurable in accordance with our assumptions stated in Section IV-A.
@@ V-B2 Tactile In-Hand Pose Estimation
To recover $\mathbf{T}_{to}$ for GrOMP, we use an in-hand pose estimation module $\mathbf{IHP}$, consisting of a convolutional variational auto-encoder [vae] and MLP, with layers as listed in Table I. These are trained simultaneously on reconstruction of the raw RGB GelSlim image (downsampled 16x, 6-channel from two fingers), minimization of KL-divergence of the latent space, and regression to ground truth object poses. This is encompassed in the loss $\mathcal{L}_{\text{ihp}}$:
@@ V-C Observation Encoding
Encoder $\mathbf{IHP}_{\text{Enc}}$ Conv2D In: $6$, Out: $12$, $3\times 3$, MaxPool($2$)
@@
Predictor $\mathbf{IHP}_{\text{MLP}}$ Linear In: $256$, Out: $84$
@@
Decoder $\mathbf{IHP}_{\text{Dec}}$ Linear In: $256$, Out: $256$
@@ V-D Task Demonstrations
We collect a tactile and proprioceptive dataset for each task the robot is to perform; both are important to our implementation of both GrOMP and IL.
@@ V-D2 Demonstration Processing
All demonstration episodes for behavior cloning a single task are sampled such that each episode contributes $T_{E}=64$ action/observation pairs, where $\mathbf{A}^{0}_{t}$ consists of $T_{a}=8$ cartesian robot poses $\mathbf{T}_{st}$ (converted to 9D pose) within the $\mathrm{SE}(3)$ robot workspace. Using position (as opposed to velocity) as an action space in this way is consistent with [chi2023] and lends itself well to GrOMP. Actions are normalized to (1,1)(-1,1) just as observations. We augment the dataset 8 times by adding noise $\sim N(0,0.1)$ to the normalized tactile shear-field. We use a train-validation split ratio of 80:20. For every 10 episodes, we then have 4096 action-observation pairs to train, and 1024 to validate. The weights of $\epsilon_{\theta}$ that produce the lowest $\mathcal{L}_{D}$ during training are selected for testing.
@@ VI Experiments and Results
We tested GrOMP as described in Section IV with DP as described in Section V, against pure vanilla DP as a baseline, with four precise assembly tasks: nut threading, peg insertion, USB insertion, and battery cover placement. We treated this as an interactive imitation learning trial where demonstration episodes are added every 10 trials to train new policies. For the DP baseline, we just tested the effect of adding episodes, beginning with 10, then 20, 40, 60, 80, and finally 100. For GrOMP we tested the same, with the effect of the projection to $\mathcal{T}$ (Eq. 4). The initial value $Q_{0}(i)$ was determined by the first 10 episodes only. We set the value update in Eq. 8 to occur every $K=1$ trial, with $\gamma=0.025$. The $\epsilon$-greedy method for selecting the projection dimensionality $i_{k}^{*}$ is performed with $\epsilon=0.1$. Each experiment—a full set of 6 different amounts of training episodes, tested 10 times each—is replicated 4 times. For each of these 4 replications, episodes are introduced in a new random order. As a result, each task is tested with 240 runs for GrOMP and 240 runs for the DP baseline. Each run is given a maximum policy horizon: $T_{E}=64$ action prediction steps to complete the task. Snapshots of the four tasks we test are found in Figs. 58.
Figure 5: Nut threading results. Top: Handoff, initialization, policy, and success snapshots. Left: GrOMP vs DP performance results as demonstrations are added to training. Right: Change of highest values in $Q(i)$ from Section IV-C over the 60-trial horizon, averaged over the 4 runs of GrOMP. Filled area surrounding curves represents $\sigma$ (the initial projection loss in Eq. 7 sometimes yielded $i_{k}^{*}=1$ but $i_{k}^{*}=2$ was the eventual result in all 4 runs of USB insertion).
@@ VI-A Nut Threading
Figure 6: Peg insertion results. Top: Handoff, initialization, policy, and success snapshots. Left: GrOMP vs DP performance results as demonstrations are added to training. Right: Change of highest values in $Q(i)$ from Section IV-C over the 60-trial horizon, averaged over the 4 runs of GrOMP. Filled area surrounding curves represents $\sigma/8$ (high variance in the result of Eq. 9 occurred between the 4 runs of peg insertion).
@@ VI-B Peg Insertion
Figure 7: USB insertion results. Top: Handoff, initialization, policy, and success snapshots. Left: GrOMP vs DP performance results as demonstrations are added to training. Right: Change of highest values in $Q(i)$ from Section IV-C over the 60-trial horizon, averaged over the 4 runs of GrOMP. Filled area surrounding curves represents $\sigma$ (Eq. 9 always yielded $i_{k}^{*}=2$ in all 4 runs of USB insertion).
@@ VI-C USB Insertion
Figure 8: Battery cover placement results. Top: Handoff, initialization, policy, and success snapshots. Left: GrOMP vs DP performance results as demonstrations are added to training. Right: Change of highest values in $Q(i)$ from Section IV-C over the 60-trial horizon, averaged over the 4 runs of GrOMP. Filled area surrounding curves represents $\sigma$ (Eq. 9 yielded $i_{k}^{*}=2$ most commonly, with $i_{k}^{*}=3$ being selected near the end of 1 of 4 runs of battery cover placement).
@@ VII Discussion
Our results suggest that GrOMP improves task performance over vanilla Diffusion Policy (DP). One key effect of GrOMP is the tendency to avoid unrecoverable states in contrast to vanilla IL, especially when the object shifts in the grasp. For example, nut threading with DP alone may exhibit behavior where the nut contacts the bolt such that it tilts heavily within the grasp. GrOMP allows the robot to position itself such that the nut remains vertical, even though this robot configuration does not occur in DP’s expert dataset.
GrOMP’s behavior is beneficially reactive to disturbances. This same kind of tilting can happen in peg insertion, USB insertion, and battery cover placement as well. A secondary recovery strategy requires obstacle avoidance. For instance, our peg insertion task is tested with a hole that does not have an excess of surface surrounding its mouth, instead there is empty space. If the peg is erroneously plunged into this empty space rather than the hole, a re-lift of the peg is required to recover, but this behavior does not appear in the expert dataset, nor is it likely to occur with GrOMP’s projection. GrOMP is more likely to prevent this kind of erroneous behavior to begin with, leading to our more successful peg insertion results seen in Fig. 6.
@@ VII-A Anomalous Baseline Performance
The reason for this paradoxical outcome is debatable, though we hypothesize some form of overfitting is occurring, possibly due to suboptimal, repetitive, or noisy demonstrations. There are also several hyperparameter and architecture choices when it comes to Diffusion Policy. We do note that testing over this set is prohibitively expensive given the need for real-world rollouts. Our results show that GrOMP successfully improves upon Diffusion Policy for these tasks, independent of the possible existence of suboptimal demonstrations or design choices.
@@ VII-B Limitations
Future improvements may be inspired by GrOMP’s limitations. GrOMP currently does not learn a temporal manifold, meaning projecting to $\mathcal{T}$ will not necessarily progress the task temporally. DP retains full responsibility for task completion, while GrOMP attempts to prevent accumulating errors. GrOMP also requires a reliable in-hand pose estimator. Any improvements in the pose estimation method will serve to improve GrOMP. For instance, we demonstrated a predictable $\mathrm{SE}(2)$ estimator given our assumptions (Sec. V-B). GrOMP can be improved if $\mathbf{IHP}$ were expanded to $\mathrm{SE}(3)$ as some works have attempted  [t3]. We note that tactile may not be available for all systems. A secondary solution would be to use vision techniques to estimate poses; however, these techniques are known to struggle with occlusion.
Our hope is that the inspiration behind GrOMP leads to the consideration of more geometry-based methods on top of the benefits that imitation learning already provides. Along this research path, the data efficiency and precision required by manipulation in fixtureless industrial assembly can be fully realized.